OpenStudy (anonymous):

Choose the correct solution of the given system of equations. y + 2x = -1 3y - x = 4

OpenStudy (anonymous):

Solutions a. (-1,1) b. (1,1) c. no solution d. none of the above

OpenStudy (anonymous):

Is this the one where you can subtract them from each other? Or sub?

OpenStudy (anonymous):

ya

OpenStudy (whpalmer4):

Do you know how to solve the system of equations?

OpenStudy (anonymous):

ya. sort of

OpenStudy (anonymous):

for the most part

OpenStudy (whpalmer4):

Okay, so solve it :-)

OpenStudy (anonymous):

:( your no fun. :) lol

OpenStudy (anonymous):

I always seem to end up with really long numbers like irrational decimals

OpenStudy (anonymous):

I sub x with ) for the first equation and got -1, but i don't know if I'm doing it right lol

OpenStudy (anonymous):

how'd u do it?

OpenStudy (anonymous):

@whpalmer4 , is it b?

OpenStudy (anonymous):

i think it might be b

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

it could be a

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

y+2(0)=-1 y+0=-1 (add 0 to both the 0 & the -1) 0 cancels out and you get y=-1

OpenStudy (anonymous):

I think it might be A, but don't lean on me lol

OpenStudy (anonymous):

You would subtract 0 from both sides In this 1 it doesnt really matter, but u would subtract

OpenStudy (anonymous):

The answer is a

OpenStudy (anonymous):

Good night

OpenStudy (anonymous):

Yea, I meant subtract...Goodnight

OpenStudy (whpalmer4):

\[y + 2x = -1\]\[3y - x = 4\] We can easily solve the first equation for \(y\), or the second equation for \(x\). Let's do the former: \[y + 2x = -1\]\[y = -1-2x\]Now substitute that into the other equation: \[3(-1-2x)-x = 4\]\[-3-6x-x = 4\]\[-7x=7\]\[x=-1\]Now put that in the substitution equation to find \(y\):\[y = -1-2(-1)\]\[y = -1+2\]\[y = 1\]So our solution is \((-1,1)\) Checking: \[1+2(-1) = -1\checkmark\]\[3(1)-(-1) = 4\checkmark\]